OpenStudy (anonymous):

The product of two positive real numbers is 20. Find the least possible sum of two such real numbers

OpenStudy (anonymous):

9 is the least possible sum the 2 nums r 4 and 5 if Im right ;)

OpenStudy (anonymous):

@sneakergyrl is right if the numbers should be distinct otherwise the numbers would be \[2\sqrt{5}\] and \[2\sqrt{5}\] And the least sum would be \[4\sqrt{5}\]

OpenStudy (anonymous):

I think you should go with 4 and 5.

OpenStudy (anonymous):

do you know how to solve it using calculus?

OpenStudy (anonymous):

thanks @ShailKumar

OpenStudy (anonymous):

So, if the numbers be x and y, then x*y = 20. We want to minimize, s = x+y So, s = x+20/x Now find ds/dx and put it equal to zero to solve for x.

OpenStudy (anonymous):

So would you have 1+20/x^2=0?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

1-20/x^2=0

OpenStudy (anonymous):

can you have a square root for your answer?

OpenStudy (anonymous):

Yes, you are getting x = 2 sqr root 5, which is a real number.

OpenStudy (anonymous):

Thank you!